A mass m hanging from a spring with spring constant k oscillates with frequency f. What change to this oscillator would make it oscillate with frequency 2f?

A.  increasing the spring constant to 2k

B.  decreasing the mass to m/4

C.  increasing the mass to 4m

D.  decreasing the spring constant to k/4
Answer #1

T = 2 \pi \sqrt(m/k)

we need to change f to 2f

which means T should be changed to T/2 (because f = 1/T)

So we can do it by decreasing mass to m/4 , increasing k to 4k